Solve: 23x/2 = 46
Solve: 23x/2 = 46
Solve: 23x/2 = 46
Solve: 23x/2 = 46
Solve: 23x/2 = 46

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

23x/2 = 46

x = (46× 2)/23

x = 2× 2

x = 4

Thanks